OpenStudy (anonymous):

The equation of a line is y=-1/2x-2. What is the equation of the line that is perpendicular to the first line and passes through the point (2, –2)? A. y = 2x – 2 B. y=1/2x - 3 C. y = 2x – 6 D. y = –2x + 2

OpenStudy (kc_kennylau):

y=mx+c m=100(given) y=100x+c substitute 2000,3000 3000=200000+c c=-197000 y=100x-197000 y=100x-197000

OpenStudy (anonymous):

y = mx + b (solving for b) 3000 = 100(2000) + b 3000 = 200000 + b 3000 - 200000 = b 197,000 = b so we have : y = 100x - 197,000 I agree with kc :)